OpenStudy (anonymous):

To build 12 computers, it takes four people 18 days. If two more people helped, how many days would it probably have taken to build all 12 computers? plz help, thank-you.

OpenStudy (anonymous):

to build 12 comp. one would take 4*18 days. to build 12 comp.,6 would take =4*18/6=12 days

OpenStudy (dumbcow):

find the rate it takes 1 person to build 1 computer 12/4 = 3 so each person built 3 computers in 18 days 3/18 = 1/6 rate is 1 computer per 6 days per person 6people then build 6 computers in 6 days ---> 12 computers in 12 days